4. Research on new juvenile delinquents revealed that 38% of them committed another crime.
a. What is the probability that of the last 100 new juvenile delinquents put on probation, 30 or more will commit another crime?
b. What is the probability that 40 or fewer of the delinquents will commit another crime?
c. What is the probability that between 30 and 40 of the delinquents will commit another crime?

Answer by edjones(8007)

c)
s=sqrt(npq)=sqrt(100*.38*.62)=4.854
m=np=100*.38=38
z=(x-m)/s=(29.5-38)/4.854=-1.751
z=(40.5-38)/4.854=0.515
The area under the normal curve between -1.751 and 0.515 is 0.657 the probability that between 30 and 40 of the delinquents will commit another crime.
.
Ed